OpenStudy (anonymous):

csc x(2sinx- sqrt 2) =0 Trigonometric function or not?

zepdrix (zepdrix):

What is it that you're asking? \[\Large\rm \csc x\left(2\sin x-\sqrt2\right)=0\]Solve for x?

OpenStudy (anonymous):

i need to figure out if it's a trigonometric identity or not

OpenStudy (anonymous):

I didn't mean function, sorry :x

zepdrix (zepdrix):

Are you sure you read the directions correctly? 0_o This sure looks like a "solve for x" problem. Hmm

OpenStudy (anonymous):

yeah, it says "decide whether the equation is a trigonometric identity"

zepdrix (zepdrix):

\(\Large\rm \csc x\) is equivalent to \(\Large\rm \dfrac{1}{\sin x}\). So that gives us\[\Large\rm \frac{1}{\sin x}\left(2\sin x-\sqrt2\right)=0\]Distributing that to the bracketed stuff gives us,\[\Large\rm 2-\frac{\sqrt2}{\sin x}=0\]When we're trying to verify whether or not something is an identity, we only alter one side and try to make it match the other side.

zepdrix (zepdrix):

So in this case, yes there are specific x values which will give us 0 on the left side, making this true, but it is certainly not an identity. >.<

zepdrix (zepdrix):

Weird stuff :U

OpenStudy (anonymous):

yeah, it is. Thank you a lot, though. That makes sense C: